# UniMoT: Unified Molecule-Text Language Model with Discrete Token Representation

### v1

@@ 3.3 Training Strategy
We connect the molecule encoder and Causal Q-Former, leveraging the pretrained MoleculeSTM molecule encoder [31]. The molecule encoder remains frozen while only the Causal Q-Former is updated. Both queries and text inputs are used, while only queries serve as input in subsequent stages. In our experiments, we utilize 16 queries. We employ three tailored objectives for the pretraining of the Causal Q-Former: Molecule-Text Contrastive Learning (MTC), Molecule-Text Matching (MTM), and Molecule-grounded Text Generation (MTG). The details of these objectives can be found in Appendix A.
@@
We connect the Causal Q-Former with subsequent blocks and use the objective defined in Equation (2). We employ the pretrained ChemFormer [17] as the generative model. Specifically, we leverage the SMILES encoder and the SMILES decoder provided by ChemFormer. The molecule codebook size is set to $K=2048$. As shown in Figure 2, we keep the molecule encoder, the SMILES encoder, and the SMILES decoder frozen, while updating the Causal Q-Former, the learnable codebook, and the adapter.
@@
We integrate the molecule tokenizer with the LLM using the unified vocabulary of molecule tokens and text tokens. We employ the LM objective defined in Equation (3) to pretrain the LLM. Pretraining involves molecule-to-text autoregression and text-to-molecule autoregression, aimed at enhancing UniMoT’s multi-modal comprehension and generation capabilities. To enhance efficiency, we train the LLM using LoRA tuning [14].
@@
UniMoT is fine-tuned on seven comprehension and generation tasks: molecular property prediction, molecule captioning, molecule-text retrieval, caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. We also utilize LoRA tuning to improve efficiency. This stage ensures UniMoT can accurately interpret and respond to human instructions, making it versatile and effective for diverse molecular tasks.
@@ 4.1 Molecule Comprehension Tasks
The goal of molecular property prediction is to forecast a molecule’s intrinsic physical and chemical properties. For the classification task, we incorporate eight binary classification datasets from MoleculeNet [53]. Models are tasked with generating a single prediction (“yes” or “no”). We compare UniMoT with the following baselines: KV-PLM [58], AttrMask [16], InfoGraph [40], MolCLR [48], GraphMVP [30], MoleculeSTM [31], and InstructMol [6]. The ROC-AUC (%) results on the MoleculeNet datasets are shown in Table 1. The performance of the regression task of molecular property prediction is provided in Appendix D. Compared to traditional graph learning methods and molecular LLMs like InstructMol [6], UniMoT demonstrates consistent improvements across the eight datasets, indicating its robust molecule comprehension abilities.
Table 1: ROC-AUC (%) of molecular property prediction task (classification) on the MoleculeNet [53] datasets. Bold indicates the best performance and underline indicates the second best performance.
@@
KV-PLM [58] 70.50 72.12 55.03 59.83 89.17 54.63 65.40 78.50
AttrMask [16] 67.79 75.00 63.57 58.05 75.44 73.76 75.44 80.28
InfoGraph [40] 64.84 76.24 62.68 59.15 76.51 72.97 70.20 77.64
MolCLR [48] 67.79 75.55 64.58 58.66 84.22 72.76 75.88 71.14
GraphMVP [30] 68.11 77.06 65.11 60.64 84.46 74.38 77.74 80.48
MoleculeSTM [31] 69.98 76.91 65.05 60.96 92.53 73.40 76.93 80.77
InstructMol (Vicuna-7B) [6] 70.00 74.67 64.29 57.80 91.48 74.62 68.90 82.30
UniMoT (Llama-2-7B) 71.37 76.43 65.78 59.79 92.89 75.97 78.49 83.69
@@
The molecule captioning task involves generating a comprehensive description of a molecule. We compare UniMoT with several baselines: MolT5 [11], MoMu [39], InstructMol [6], MolCA [33], and 3D-MoLM [25]. BLEU [37], ROUGE [27], and METEOR [1] are adopted as evaluation metrics. UniMoT is evaluated for molecule captioning on the PubChem [18] and CheBI-20 [11] datasets. Performance on the PubChem dataset is shown in Table 2, while the performance on the CheBI-20 dataset and some concrete examples are presented in Appendix D.
From Table 2, we observe that UniMoT consistently outperforms the baselines by a significant margin on the PubChem [18] dataset. This task is more complex than classification or regression, providing a robust measure of the model’s molecule comprehension abilities. Notably, our proposed tokenizer-based architecture surpasses the projection-based architecture (such as InstructMol [6]), Q-Former-based architecture (such as MolCA [33] and 3D-MoLM [25]), and models trained with contrastive learning strategies (such as MoMu [39]). The results demonstrate that the molecule tokenizer can generate molecule tokens with high-level molecular and textual information.
Table 2: Performance (%) of molecule captioning task on the PubChem [18] dataset. Bold indicates the best performance and underline indicates the second best performance.
[…]
MolT5-Small (T5-Small) [11] 22.5 15.2 30.4 13.5 20.3 24.0
MolT5-Base (T5-Base) [11] 24.5 16.6 32.2 14.0 21.4 26.1
MolT5-Large (T5-Large) [11] 25.9 17.3 34.1 16.4 23.4 28.0
MoMu-Small (T5-Small) [39] 22.9 16.0 31.0 13.7 20.8 24.4
MoMu-Base (T5-Base) [39] 24.7 16.8 32.5 14.6 22.1 27.2
MoMu-Large (T5-Large) [39] 26.3 18.0 34.8 16.9 24.8 28.7
InstructMol (Vicuna-7B) [6] 18.9 11.7 27.3 11.8 17.8 21.3
MolCA (OPT-125M) [33] 25.9 17.5 34.4 16.6 23.9 28.5
MolCA (OPT-1.3B) [33] 28.6 21.3 36.2 21.4 29.7 32.6
3D-MoLM (Llama-2-7B) [25] 30.3 22.5 36.8 22.3 31.2 33.1
UniMoT (Llama-2-7B) 31.3 23.8 37.5 23.7 33.6 34.8
Molecule-Text Retrieval Task.
The molecule-text retrieval task involves using a molecule to retrieve text (M2T) and using text to retrieve a molecule (T2M). We compare UniMoT with several baselines: Sci-BERT [2], KV-PLM [58], MoMu [39], MoleculeSTM [31], MolCA [33], and 3D-MoLM [25]. We report the performance of retrieval using a batch of 64 random samples and the entire test set, evaluated with the metrics of Accuracy and Recall@20. We use the checkpoint from Stage-1 of pretraining. UniMoT is evaluated on the datasets of PubChem [18], PCdes [58], and MoMu [39]. Performance on the PubChem dataset is shown in Table 3, while performances on the PCdes and MoMu datasets are presented in Appendix D. UniMoT can understand complex molecule-text interactions through the introduction of the Causal Q-Former. From Table 3, UniMoT demonstrates superior performance over the baselines on molecule-to-text retrieval. This underscores UniMoT’s capability in learning fine-grained alignment between molecules and text.
Table 3: Performance (%) of molecule-text retrieval task on the PubChem [18] dataset. Bold indicates the best performance and underline indicates the second best performance.
Model Retrieval in batch Retrieval in test set
M2T (%) T2M (%) M2T (%) T2M (%)
[…]
Sci-BERT [2] 85.3 98.7 84.2 98.4 41.7 87.3 40.2 86.8
KV-PLM [58] 86.1 98.6 85.2 98.5 42.8 88.5 41.7 87.8
MoMu (Sci-BERT) [39] 87.6 99.2 86.4 99.4 47.3 90.8 48.1 89.9
MoMu (KV-PLM) [39] 88.2 99.4 87.3 99.4 48.5 91.6 49.5 90.7
MoleculeSTM [31] 90.5 99.6 88.6 99.5 52.7 92.9 53.2 92.5
MolCA (OPT-1.3B) [33] 92.6 99.8 91.3 99.5 67.9 94.4 68.6 93.3
3D-MoLM (Llama-2-7B) [25] 93.5 100.0 92.9 99.6 69.1 95.9 70.1 94.9
UniMoT (Llama-2-7B) 93.6 100.0 92.7 99.4 69.5 96.3 69.8 94.4
@@ 4.2 Molecule Generation Tasks
Table 4: Performance of molecule generation tasks on the Mol-Instructions [12] datasets, including caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. Bold indicates the best performance, and underline indicates the second best performance.
@@
Llama [44] 0.000 0.003 59.864 0.005 0.000 0.000 0.003
Vicuna [7] 0.000 0.006 60.356 0.006 0.001 0.000 0.001
Mol-Instructions [12] 0.002 0.345 41.367 0.231 0.412 0.147 1.000
MolT5 [11] 0.112 0.546 38.276 0.400 0.538 0.295 0.773
@@
Llama [44] 0.000 0.003 28.040 0.037 0.001 0.001 0.001
Vicuna [7] 0.000 0.010 27.948 0.038 0.002 0.001 0.007
Mol-Instructions [12] 0.044 0.224 23.167 0.237 0.364 0.213 1.000
InstructMol [6] 0.129 0.610 19.664 0.444 0.539 0.400 1.000
@@
Forward Reaction Prediction
Llama [44] 0.000 0.020 42.002 0.001 0.002 0.001 0.039
Vicuna [7] 0.000 0.057 41.690 0.007 0.016 0.006 0.059
Mol-Instructions [12] 0.045 0.654 27.262 0.313 0.509 0.262 1.000
InstructMol [6] 0.536 0.967 10.851 0.776 0.878 0.741 1.000
UniMoT 0.611 0.980 8.297 0.836 0.911 0.807 1.000
Retrosynthesis
Llama [44] 0.000 0.036 46.844 0.018 0.029 0.017 0.010
Vicuna [7] 0.000 0.057 46.877 0.025 0.030 0.021 0.017
Mol-Instructions [12] 0.009 0.705 31.227 0.283 0.487 0.230 1.000
InstructMol [6] 0.407 0.941 13.967 0.753 0.852 0.714 1.000
UniMoT 0.478 0.974 11.634 0.810 0.909 0.771 1.000
We employ molecule generation tasks, which encompass caption-guided molecule generation [12], reagent prediction [12], forward reaction prediction [12], and retrosynthesis [12]. Caption-guided molecule generation involves generating molecular structures based on textual descriptions. Reagent prediction entails determining suitable reagents given reactants and products. Forward reaction prediction involves predicting probable products given specific reactants and reagents. Retrosynthesis involves deconstructing a target molecule into simpler starting materials. We compare UniMoT with the following baselines: Llama [44], Vicuna [7], Mol-Instructions [12], and InstructMol [6]. The metrics used to evaluate molecule generation tasks include Exact Match, BLEU [37], Levenshtein Distance [22], RDKit Fingerprint Similarity [20], MACCS Fingerprint Similarity [10], and Morgan Fingerprint Similarity [36]. These metrics evaluate structural similarity between generated and target molecules, along with Validity [19], which assesses the proportion of chemically valid molecules generated. We utilize the Mol-Instructions [12] datasets to evaluate the generation capabilities of UniMoT, and the results are presented in Table 4.
As the baselines generate SMILES strings and then convert them to molecules, UniMoT directly leverages the generated molecule tokens and obtains their embeddings from the learned codebook. These embeddings can be decoded to desired molecules through the pretrained adapter and SMILES decoder. Regarding the results in Table 4, UniMoT exhibits the capability to generate valid molecules with a higher degree of similarity to the target molecules compared to the baselines. UniMoT can generate molecules as if they were text, demonstrating strong generation capabilities and providing a new perspective to molecule generation tasks.
@@ 4.3 Ablation Studies
Table 5: Ablation study on the projector and representation form for the molecule captioning task using the PubChem [18] dataset.
[…]
@@
We conducted an ablation study on the cross-modal projector, with the results on the molecule captioning task shown in Table 5. The linear projection demonstrated the worst performance, indicating that the molecule features lack textual information, thus hindering effective molecule-text interactions and alignment. Additionally, we compared the performance of a Q-Former with bidirectional self-attention to a Causal Q-Former with causal self-attention. The results show that queries with causal dependency outperform those with bidirectional dependency. This demonstrates that input with left-to-right causal dependency aligns with the unidirectional attention mechanism in LLMs, leading to improved performance.
@@
We compare the performance of continuous causal embeddings and discrete tokens quantized from causal embeddings as inputs to LLMs. As shown in Table 5, continuous embeddings demonstrate better performance than discrete tokens in understanding molecules. This result is reasonable since the quantization process causes information loss in discrete tokens. However, we still use discrete token representation to facilitate the autoregressive training paradigm of LLMs, which supports the unification of comprehension and generation tasks. To achieve this unification, we unavoidably sacrifice some performance in comprehension tasks.
Model Size and Tuning Stategy.
Table 6: Ablation study on the model size and tuning strategy for the molecule captioning task using the PubChem [18] dataset.
[…]
Llama-2-7B LoRA Tuning 31.3 23.8 37.5 23.7 33.6 34.8
Llama-2-7B Full Tuning 32.0 24.6 38.3 24.3 34.7 35.6
Llama-2-13B LoRA Tuning 31.8 24.3 38.0 24.1 34.4 35.3
We conducted a comparison of molecule captioning performance across various model sizes and tuning strategies, as illustrated in Table 6. Our findings indicate that scaling up the LLM to 13B or adopting a full tuning strategy yields only marginal improvements in performance compared to using Llama-2-7B with LoRA tuning. While larger models and full tuning strategies might offer slight gains in performance, they come at a significant cost in terms of efficiency. Considering the trade-off between achieving high performance and maintaining efficiency, we have chosen to utilize Llama-2-7B with LoRA tuning in our experiments. This ensures that our model remains both powerful and practical.
@@ 5 Conclusion
This work introduces UniMoT, an innovative advancement in the field of molecular-textual understanding and generation, which successfully unifies these two distinct modalities under a coherent framework. By introducing a Vector Quantization-driven tokenizer, UniMoT overcomes previous architectural limitations where molecule and text modalities were not treated equally. This molecule tokenizer transforms molecules into sequences of discrete tokens, embedding high-level molecular and textual information. The LLM vocabulary is expanded with molecule tokens mapped from the learned codebook. Moreover, by employing a four-stage training scheme, UniMoT has emerged as a versatile multi-modal LLM, adept at handling both molecule-to-text and text-to-molecule tasks. Extensive empirical evaluations demonstrate that UniMoT attains state-of-the-art performance across diverse molecule comprehension and generation tasks.

### v2

@@ 3.3 Training Strategy
We connect the molecule encoder and Causal Q-Former, leveraging the pretrained MoleculeSTM molecule encoder Liu et al. (2023a). The molecule encoder remains frozen while only the Causal Q-Former is updated. Both queries and text inputs are used, while only queries serve as input in subsequent stages. In our experiments, we utilize 16 queries. We employ three tailored objectives for the pretraining of the Causal Q-Former: Molecule-Text Contrastive Learning (MTC), Molecule-Text Matching (MTM), and Molecule-grounded Text Generation (MTG). The details of these objectives can be found in Appendix A.
@@
We connect the Causal Q-Former with subsequent blocks and use the objective defined in Equation (2). We employ the pretrained ChemFormer Irwin et al. (2022) as the generative model. Specifically, we leverage the SMILES encoder and the SMILES decoder provided by ChemFormer. The molecule codebook size is set to $K=2048$. As shown in Figure 2, we keep the molecule encoder, the SMILES encoder, and the SMILES decoder frozen, while updating the Causal Q-Former, the learnable codebook, and the adapter.
@@
We integrate the molecule tokenizer with the LLM using the unified vocabulary of molecule tokens and text tokens. We employ the LM objective defined in Equation (3) to pretrain the LLM. Pretraining involves molecule-to-text autoregression and text-to-molecule autoregression, aimed at enhancing UniMoT’s multi-modal comprehension and generation capabilities. To enhance efficiency, we train the LLM using low-rank adaptation (LoRA) Hu et al. (2021).
@@
UniMoT is fine-tuned on seven comprehension and generation tasks: molecular property prediction, molecule captioning, molecule-text retrieval, caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. We also utilize LoRA to improve efficiency. This stage ensures UniMoT can accurately interpret and respond to human instructions, making it versatile and effective for diverse molecular tasks.
@@ 4.1 Molecule Comprehension Tasks
The goal of molecular property prediction is to forecast a molecule’s intrinsic physical and chemical properties. For the classification task, we incorporate eight binary classification datasets from MoleculeNet Wu et al. (2018). Models are tasked with generating a single prediction (“yes” or “no”). We compare UniMoT with the following baselines: KV-PLM Zeng et al. (2022), AttrMask Hu et al. (2019), InfoGraph Sun et al. (2019), MolCLR Wang et al. (2021), GraphMVP Liu et al. (2019), MoleculeSTM Liu et al. (2023a), and InstructMol Cao et al. (2023). The ROC-AUC (%) results on the MoleculeNet datasets are shown in Table 1. The performance of the regression task of molecular property prediction is provided in Appendix D. Compared to traditional graph learning methods and molecular LLMs like InstructMol Cao et al. (2023), UniMoT demonstrates consistent improvements across the eight datasets, indicating its robust molecule comprehension abilities.
@@
MolT5-Small (T5-Small) 22.5 15.2 30.4 13.5 20.3 24.0
MolT5-Base (T5-Base) 24.5 16.6 32.2 14.0 21.4 26.1
MolT5-Large (T5-Large) 25.9 17.3 34.1 16.4 23.4 28.0
MoMu-Small (T5-Small) 22.9 16.0 31.0 13.7 20.8 24.4
MoMu-Base (T5-Base) 24.7 16.8 32.5 14.6 22.1 27.2
MoMu-Large (T5-Large) 26.3 18.0 34.8 16.9 24.8 28.7
InstructMol (Vicuna-7B) 18.9 11.7 27.3 11.8 17.8 21.3
MolCA (OPT-125M) 25.9 17.5 34.4 16.6 23.9 28.5
MolCA (OPT-1.3B) 28.6 21.3 36.2 21.4 29.7 32.6
3D-MoLM (Llama-2-7B) 30.3 22.5 36.8 22.3 31.2 33.1
UniMoT (Llama-2-7B) 31.3 23.8 37.5 23.7 33.6 34.8
Table 2: Performance (%) of molecule captioning task on the PubChem Kim et al. (2023) dataset. Bold indicates the best performance and underline indicates the second best performance.
@@
The molecule captioning task involves generating a comprehensive description of a molecule. We compare UniMoT with several baselines: MolT5 Edwards et al. (2022), MoMu Su et al. (2022), InstructMol Cao et al. (2023), MolCA Liu et al. (2023b), and 3D-MoLM Li et al. (2024). BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005) are adopted as evaluation metrics. UniMoT is evaluated for molecule captioning on the PubChem Kim et al. (2023) and ChEBI-20 Edwards et al. (2022) datasets. Performance on the PubChem dataset is shown in Table 2, while the performance on the ChEBI-20 dataset and some concrete examples are presented in Appendix D. The ChEBI-20 dataset replaces molecular names with “the molecule” to focus on properties. However, predicting molecular names reflects the model’s structural understanding, so we conducted the main experiments on PubChem.
From Table 2, we observe that UniMoT consistently outperforms the baselines by a significant margin on the PubChem Kim et al. (2023) dataset. This task is more complex than classification or regression, providing a robust measure of the model’s molecule comprehension abilities. Notably, our proposed tokenizer-based architecture surpasses the projection-based architecture (such as InstructMol Cao et al. (2023)), Q-Former-based architecture (such as MolCA Liu et al. (2023b) and 3D-MoLM Li et al. (2024)), and models trained with contrastive learning strategies (such as MoMu Su et al. (2022)). This demonstrates that the tokenizer-based architecture achieves better molecule-text alignment through autoregressive molecule-to-text and text-to-molecule pretraining compared to other architectures. Details and More Results of Experiments can be found in Appendix D.
@@ 4.2 Molecule Generation Tasks
We employ molecule generation tasks, which encompass caption-guided molecule generation Fang et al. (2023), reagent prediction Fang et al. (2023), forward reaction prediction Fang et al. (2023), and retrosynthesis Fang et al. (2023). Caption-guided molecule generation involves generating molecular structures based on textual descriptions. Reagent prediction entails determining suitable reagents given reactants and products. Forward reaction prediction involves predicting probable products given specific reactants and reagents. Retrosynthesis involves deconstructing a target molecule into simpler starting materials. We compare UniMoT with the following baselines: Llama Touvron et al. (2023), Vicuna Chiang et al. (2023), Mol-Instructions Fang et al. (2023), and InstructMol Cao et al. (2023). The metrics used to evaluate molecule generation tasks include Exact Match, BLEU Papineni et al. (2002), Levenshtein Distance Levenshtein and others (1966), RDKit Fingerprint Similarity Landrum and others (2006), MACCS Fingerprint Similarity Durant et al. (2002), and Morgan Fingerprint Similarity Morgan (1965). These metrics evaluate structural similarity between generated and target molecules, along with Validity Kusner et al. (2017), which assesses the proportion of chemically valid molecules generated.
We utilize the Mol-Instructions Fang et al. (2023) benchmark to evaluate the generation capabilities of UniMoT. The results of caption-guided molecule generation and reagent prediction are presented in Table 3, and the results of other tasks are in Appendix D. The caption-guided molecule generation task, the reverse of molecule captioning, is conducted using the PubChem Kim et al. (2023) dataset, while the other tasks utilize the USPTO Fang et al. (2023) dataset. As the baselines generate SMILES strings and then convert them to molecules, UniMoT directly leverages the generated molecule tokens and obtains their embeddings from the learned codebook. These embeddings can be decoded to desired molecules through the pretrained adapter and SMILES decoder. As shown in Table 3, UniMoT generates valid molecules with a higher degree of similarity to the target molecules compared to the baselines. This is because UniMoT can generate molecules as if they were text, which is fundamentally different from adapter-based architectures. UniMoT demonstrates strong generation capabilities and offers a new perspective on these tasks.
@@
Llama 0.000 0.003 59.864 0.005 0.000 0.000 0.003
Vicuna 0.000 0.006 60.356 0.006 0.001 0.000 0.001
Mol-Instructions 0.002 0.345 41.367 0.231 0.412 0.147 1.000
MolT5 0.112 0.546 38.276 0.400 0.538 0.295 0.773
@@
Llama 0.000 0.003 28.040 0.037 0.001 0.001 0.001
Vicuna 0.000 0.010 27.948 0.038 0.002 0.001 0.007
Mol-Instructions 0.044 0.224 23.167 0.237 0.364 0.213 1.000
InstructMol 0.129 0.610 19.664 0.444 0.539 0.400 1.000
@@
Table 3: Performance of molecule generation tasks on the Mol-Instructions Fang et al. (2023) benchmark, including caption-guided molecule generation and reagent prediction. Bold indicates the best performance, and underline indicates the second best performance.
@@ 4.3 Ablation Studies
Projector Input to LLM BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
@@
(a) Ablation study on the projector and representation form.
Architecture Codebook Size BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
Llama-2-7B 512 28.7 20.5 33.2 20.7 29.6 30.2
Llama-2-7B 1024 29.5 21.3 34.5 21.8 30.9 31.5
Llama-2-7B 2048 31.3 23.8 37.5 23.7 33.6 34.8
Llama-2-7B 4096 31.1 23.6 37.1 23.5 33.2 34.3
(b) Ablation study on the codebook size.
Table 4: Ablation studies on the molecule captioning task using the PubChem dataset.
We conducted an ablation study on the cross-modal projector, with the results on the molecule captioning task shown in Table 4(a). The linear projection demonstrated the worst performance, indicating that the molecule features lack textual information, thus hindering effective molecule-text interactions and alignment. Additionally, we compared the performance of a Q-Former with bidirectional self-attention to a Causal Q-Former with causal self-attention in the second and third rows. The results show that queries with causal dependency outperform those with bidirectional dependency. This demonstrates that input with left-to-right causal dependency aligns with the unidirectional attention mechanism in LLMs, leading to improved performance.
@@
We compared the performance of continuous causal embeddings and discrete tokens, quantized from causal embeddings, as inputs to LLMs in the third and fourth rows of Table 4(a). Continuous embeddings demonstrate better performance than discrete tokens in understanding molecules. This result is reasonable since the quantization process causes information loss in discrete tokens. However, we still use discrete token representation to facilitate the autoregressive training paradigm of LLMs, which supports the unification of comprehension and generation tasks. To achieve this unification, we unavoidably sacrifice some performance in comprehension tasks.
Codebook Size.
We conducted experiments with different molecule codebook sizes and reported the performance on the molecule captioning task. The performance is shown in Table 4(b). The results demonstrate that the codebook size of 2048 consistently provides the best performance. This choice balances model complexity and performance. A larger codebook could capture more subtle interactions between molecules and text. However, there may be some codes that are not often used. A smaller codebook may result in nearby embeddings being assigned the same code, which reduces the granularity of the representation. More ablation studies are presented in Appendix E.
@@ 5 Conclusion
This work introduces UniMoT, a framework that unifies the modalities of molecules and text. By adopting a tokenizer-based architecture, UniMoT addresses previous limitations where the molecule and text modalities are not treated equally. The molecule tokenizer converts molecules into sequences of discrete tokens, embedding high-level molecular and textual information. The LLM vocabulary is expanded with molecule tokens mapped from a learned codebook. Through a four-stage training scheme, UniMoT has become a versatile multi-modal LLM, capable of handling both molecule-to-text and text-to-molecule tasks. Extensive empirical evaluations show that UniMoT achieves state-of-the-art performance across diverse molecule comprehension and generation tasks.
Although UniMoT excels in molecule-to-text and text-to-molecule tasks, it has yet to be extensively tested on more complex tasks like molecule editing, which require precise structural modifications. Additionally, limited annotated data in the molecular domain restricts UniMoT’s training, hindering its ability to fully learn and generalize molecular structures and properties. To improve its effectiveness, addressing data scarcity is crucial. Furthermore, expanding evaluations to include a broader range of real-world scenarios will offer a more comprehensive understanding of the model’s robustness and generalizability.
Acknowledgments
Q. Yao is supported by National Natural Science Foundation of China (under Grant No.92270106), Beijing Natural Science Foundation (under Grant No.4242039), and CCF-Zhipu Large Model Innovation Fund (CCF-Zhipu202402)